The difference between two numbers is 30 and the larger number is eight fifths of the smaller number. Find the two numbers.

Answer by macston(5194)

.
L=larger number=(8/5)S ; S=smaller number
.
L-S=30 Substitute for L.
(8/5)S-(5/5)S=30
(3/5)S=30 Multiply by 5/3.
S=50 ANSWER 1: The smaller number is 50.
L=(8/5)50=400/5=80 ANSWER 2: The larger number is 80.